Solve the following system by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
10x+2y=7
y=-5x+3

Answer by stanbon(75887)

Solve the following system by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
10x+2y=7
y=-5x+3
----------
Substitute to get:
10x+2(-5x+3)=7
10x-10x+6 = 7
6=7
This is a contradiction that arose because we assumed the
"x" and the "y" variables in the two equation meant the
same number.
That assumption led to the contradiction that 6=7.
Our assumption was wrong.
The system is said to be "inconsistent".
=============
Cheers,
Stan H.